Open Access publishing at Dept of Information Technology,

Uppsala University

Kurt Otto 2014-12-05

OA policy for publications with IT copyright

Doctoral theses and Licentiate theses from the IT department are published in the open archive DiVA (Digitala Vetenskapliga Arkivet).

Technical reports from the IT department are openly archived at

http://www.it.uu.se/research/publications/reports/

The technical reports (TR) from the division of Scientific Computing are also in DiVA with links to full texts in the TR archive.

\underline{Why DiVA?}

That the IT department should be well represented in DiVA with good order among the items serves several purposes:

1. To provide a reliable and reasonably complete database of our publications for research colleagues

2. Convenient generation of publication lists on our web pages

3. To ascertain appropriate faculty funding, which will partly be determined by the number of refereed publications in DiVA

4. To be a source for bibliometric studies

OA policy for publications without IT copyright

There are (at least) two adequate ways for IT to publish Open Access:

1. Publish in a subscription-based journal and make a copy available in an open archive, which conveniently means DiVA

2. Publish in an open access journal As far as possible, avoid the expensive way:

• Publish in a hybrid journal with open access option

Division of Scientific Computing in DiVA

The publication count (2014-12-04) as of 1999:

• 605 refereed items

• 48 doctoral theses

• 40 licentiate theses

• 237 technical reports

\underline{Challenges}

• To get the authors to upload full texts in DiVA • Dealing with OA reluctant journals

DiVA extraction tour

Let’s go to http://www.it.uu.se/research/scicomp/publications